# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

### v1

@@ 2 Preliminaries
In addition, recall that the $w^{*}$-topology on $E^{\prime}$ is the smallest topology, such that all maps $\langle\,\cdot\,,x\rangle_{E^{\prime},E}:E^{\prime}\to\mathbb{R}$ are continuous. The $w^{*}$-topology has some highly useful properties, which are summarized in the following result, whose first two assertions can be found in in [32, Theorem 2.6.18 and Theorem 2.6.23] or [11, Theorem V.3.1 and Theorem V.5.1]. The last assertion follows from the fact that compact metric spaces are separable.
@@
Note that for the proofs of our main results we need the existence of such sets $\mathfrak{D}$, and another key result, namely Proposition 4.1, also need such sets in its formulation. In addition, the fact that the $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable is often used in the the proofs as it allows to deal with sequences rather than nets and this in turn opens the door for limit theorems of integrals.
In this work, we need two $\sigma$-algebras on $E$: The first one is the usual Borel $\sigma$-algebra $\mathfrak{B}(E)$ that is generated by the norm topology on $E$ and the second one is $\sigma(E^{\prime})$, that is, the smallest $\sigma$-algebra on $E$, for which all bounded, linear $x^{\prime}:E\to\mathbb{R}$ are measurable, where, as usual, $\mathbb{R}$ is equipped with $\mathfrak{B}:=\mathfrak{B}(\mathbb{R})$. Similarly, we write $\mathfrak{B}^{d}:=\mathfrak{B}(\mathbb{R}^{d})$. Throughout this work, we assume that $\mathbb{R}^{d}$ is equipped with $\mathfrak{B}^{d}$.
In general, we have $\sigma(E^{\prime})\subset\mathfrak{B}(E)$ and the following theorem, which is a consequence of e.g. [10, Theorem E.9] and which can be also found in [21, Proposition 1.1.1], shows that for separable $E$ both $\sigma$-algebras coincide.
@@
Let $(\Omega,{\cal A})$ be a measurable space, $E$ be a Banach space, and $X:\Omega\to E$. Then $X$ is called weakly measurable if it is $({\cal A},\sigma(E^{\prime}))$-measurable. By a standard measurability argument, see e.g. [27, Corollary 1.82], we see that this is equivalent to the $({\cal A},\mathfrak{B})$-measurability of $\langle x^{\prime},X\rangle_{E^{\prime},E}:\Omega\to\mathbb{R}$ for all $x^{\prime}\in E^{\prime}$. Moreover, $X$ is called strongly measurable if it is $({\cal A},\mathfrak{B}(E))$-measurable and $f(\Omega)$ is separable. If $E$ is separable, then one can quickly check with Theorem 2.2 that weak and strong measurability are equivalent. Finally, if $(\Omega,{\cal A},\mathrm{P})$ is a probability space, then a strongly measurable $X:\Omega\to E$ is called a random variable, and in this case its image measure $\mathrm{P}_{X}$ on $\mathfrak{B}(E)$ is called the distribution of $X$. Of course, if $E$ is separable, every weakly measurable $X:\Omega\to E$ is a random variable, and for this reason we sometimes simply speak of measurable $X$.
@@
It turns out, that in this case the limit $\mathrm{P}$ is unique, see e.g. [46, Chapter I.3.7]. Of course, for $E=\mathbb{R}^{d}$, Lévy’s continuity theorem shows that weak convergence of probability measures can be characterized by the pointwise convergence of the corresponding characteristic functions, see also Theorem C.3. A generalization to the case of separable Banach spaces $E$ is recalled in Theorem C.6. Finally, like in the finite dimensional case, we say that a sequence $(X_{n})$ of $E$-valued random variables converge to an $E$-valued random variable $X$ in distribution, if the distributions of $(X_{n})$ weakly converge to the distribution of $X$.
@@
$\mathbb{E}X:=\int_{\Omega}X\,\mathrm{d}\mathrm{P}\,.$
We refer to Appendix A, which collects some information on Bochner integrals. Moreover, for $X\in{{\cal L}_{2}(\mathrm{P},E)}$ and $Y\in{{\cal L}_{2}(\mathrm{P},E)}$, where $F$ is another Banach space, the cross covariance of $X$ and $Y$ is the bounded linear operator $\operatorname{cov}(X,Y):F^{\prime}\to E$ defined by the Bochner integrals
@@
Recall that $\operatorname{cov}(X,Y):F^{\prime}\to E$ is a bounded linear operator, which is nuclear and thus compact, see e.g. [46, Theorem 2.5 in Chapter III.2] and Lemma B.7 for an alternative argument. In particular, the covariance of $X$, which is defined by
@@
where $X_{i}$ and $Y_{j}$ denote the $i$th, respectively $j$th, component function of $X$ and $Y$. Note that in this description, we formally write elements of $(\mathbb{R}^{n})^{\prime}$ as row vectors and identify $(\mathbb{R}^{n})^{\prime}$ by $\mathbb{R}^{n}$ via the canonical isomorphism given by the transposition, that is $\,\cdot\,^{\mathsf{T}}:(\mathbb{R}^{n})^{\prime}\to\mathbb{R}^{n}$. In other words $\operatorname{cov}(X,Y)(y^{\prime})=K_{X,Y}\cdot(y^{\prime})^{\mathsf{T}}$ for all $y^{\prime}\in(\mathbb{R}^{n})^{\prime}$.
@@
Moreover, a probability measure $\mathrm{Q}$ on $\sigma(E^{\prime})=\mathfrak{B}(E)$ is a Gaussian measure, if the image measure $\mathrm{Q}_{x^{\prime}}$ on $(\mathbb{R},\mathfrak{B})$ is Gaussian measure for all $x^{\prime}\in E^{\prime}$.
It turns out, that it suffices to verify the Gaussianity for a $\mathfrak{D}\subset B_{E^{\prime}}$ that is $w^{*}$-dense in $B_{E^{\prime}}$, see Theorem 5.1. In addition, it directly follows from the definition that $\mathrm{Q}$ is a Gaussian measure on $\sigma(E^{\prime})$, if and only if the identity map $\operatorname{id}_{E}:E\to E$ defined on the probability space $(E,\sigma(E^{\prime}),\mathrm{Q})$ is a Gaussian random variable. Finally, every affine linear transformation of a Gaussian random variable is a Gaussian random variable. To be more precise, if $X:\Omega\to E$ is a Gaussian random variable, $T:E\to F$ is a bounded linear operator mapping into a separable Banach space $F$, and $y\in F$, then $y+T\circ X$ is Gaussian random variable.
Recall that Fernique’s theorem, see e.g. [30, Corollary 3.2], shows that Gaussian random variables $X:\Omega\to E$ on some probability space $(\Omega,{\cal A},\mathrm{P})$ satisfy
@@
Like in the finite dimensional case, the expectation and covariance uniquely determine a Gaussian measure. Namely, if $\mathrm{P}$ and $\mathrm{Q}$ are Gaussian measures on $\sigma(E^{\prime})$ with $\mu_{\mathrm{P}}=\mu_{\mathrm{Q}}$ and $\operatorname{cov}(\mathrm{P})=\operatorname{cov}(\mathrm{Q})$, then we have $\mathrm{P}=\mathrm{Q}$. This can be quickly verified with the help of characteristic functions, e.g. by combining [46, Theorem 2.2 in Chapter IV.2.1] with [46, Proposition 2.8 in Chapter IV.2.4], where the latter is restated in a simplified form in Proposition C.5. For this reason, we write
@@
for Gaussian measures $\mathrm{Q}$. More information on characteristic functions of Gaussian measures and random variables are collected in Appendix C. Finally, we call
@@
the Gaussian Hilbert space associated to $X$, where, as usual, we identify the functions $\langle x^{\prime},X-\mathbb{E}X\rangle_{E^{\prime},E}$ with their $\mathrm{P}$-equivalence classes. It turns out that ${\cal G}(X)$ is indeed a Gaussian Hilbert space in the sense of [23]. For this and additional properties of ${\cal G}(X)$ we refer to Lemma 5.4.
@@
of $\mathrm{Q}$ given $F$. We write $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):=\mathrm{Q}(\,\cdot\,|\,\cdot\,)$ and speak of the regular conditional probability of $X$ given $Y$. Note that this is justified as $\mathrm{P}(B|y)$ describes the probability of $X\in B$ under the condition that we have observed $y=Y(\omega)$. Consequently, our goal is to describe $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ modulo $\mathrm{P}_{Y}$-sets. This task will be accomplished by
@@
In the finite dimensional case, the solution of this task is folklore, see e.g. [36, Appendix A.2], [5, Chapter 2.3], [19, Lemma 4.3], [41, Theorem 6.20], and also [35, Section 4]. To formulate the solution in the most general setting, let us recall that a $n\times m$-Matrix $B$ is said to be a Moore-Penrose pseudo-inverse of a given $m\times n$-matrix $A$, if the following four equations hold true:
@@
It can be shown that for every $m\times n$-matrix $A$ there exists exactly one Moore-Penrose pseudo-inverse, which we denote by $A^{\dagger}$, see e.g. [34, Theorem 1]. Similarly, if we have a linear map $A:\mathbb{R}^{n}\to\mathbb{R}^{m}$, then a linear map $B:\mathbb{R}^{m}\to\mathbb{R}^{n}$ is a Moore-Penrose pseudo-inverse of $A$, if the four equations above are satisfied. The uniqueness of such a pseudo-inverse can be shown analogously to the matrix case and the existence follows from the matrix case by using the matrices representing $A$ and $B$. We again write $A^{\dagger}$ for the pseudo-inverse.
@@
Finally, we also need the concept of conditional expectations, which, at least for $\mathbb{R}$-valued random variables, is probably better known than regular conditional probabilities. Its generalization to $E$-valued random variables is straightforward. Namely, if $E$ is a Banach space, $(\Omega,{\cal A},\mathrm{P})$ is a probability space, ${\cal C}\subset{\cal A}$ is a sub-$\sigma$-algebra, and $X\in{{\cal L}_{1}(\mathrm{P},E)}$, then a $\mathrm{P}$-integrable and strongly ${\cal C}$-measurable $Z:\Omega\to E$ is called a conditional expectation of $X$ given ${\cal C}$, if
@@
Like in the $\mathbb{R}$-valued case, the conditional expectation of $X$ given ${\cal C}$ is $\mathrm{P}$-almost surely unique, and hence we write $\mathbb{E}(X|{\cal C}):=Z$ and say that $Z$ is a version of $\mathbb{E}(X|{\cal C})$. Moreover, if ${\cal C}$ is generated by a random variable $Y$, that is ${\cal C}=\sigma(Y)$, we write $\mathbb{E}(X|Y):=\mathbb{E}(X|\sigma(Y))=\mathbb{E}(X|{\cal C})$. Banach space valued conditional expectations enjoy a lot of the properties of $\mathbb{R}$-valued conditional expectations, see Appendix D. In addition, Appendix D also contains a detailed description of the relationship between regular conditional probabilities and conditional expectations in the Banach space setting.
@@ 3 Main Results
As mentioned in the introduction, [6, Corollary 3.10.3] has already shown that there is a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, such that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian for all $y$. Later, in Theorem 3.5 we will recover such a version, too. Moreover, for more specific cases, formulas for the mean and covariance have also been derived by other authors as laid out in the introduction. However, we could not find such formulas in the full generality presented in Theorem 3.1. Since these formulas play a crucial role for our approximation scheme in the case of infinite dimensional $Y$, we thus decided to present Theorem 3.1 as a main result.
Our next goal is to generalize Theorem 3.1 to infinite dimensional $Y$. Here we recall that we cannot expect a literal generalization. To illustrate this, let us assume for simplicity that $H:=F$ is an infinite dimensional Hilbert space. Using the canonical Riesz-Fréchet-identification of $H^{\prime}$ with $H$, we can then view $\operatorname{cov}(Y)$ as a compact linear operator $\operatorname{cov}(Y):H\to H$. Now, [18, Theorem 2.4] shows that a bounded linear operator $A:H\to H$ has a Moore-Penrose inverse if and only if $\operatorname{ran}A$ is closed. However, it is well-known that compact operators have a closed range, if and only if their range is actually finite dimensional, see e.g. [32, Proposition 3.4.6]. Consequently, the Moore-Penrose inverse $(\operatorname{cov}Y)^{\dagger}$ exists if and only if $\operatorname{ran}\operatorname{cov}(Y)$ is finite dimensional. The latter is only possible, if the range of $Y$ is almost surely contained in a finite-dimensional subspace of $F$, see e.g. [29, Theorem 2.1] or [42, Theorem 2.43].
In the following, we will resolve this issue by considering a sequence $(Y_{n})$ of suitably chosen finite dimensional projections of $Y$. For each $Y_{n}$ we can then apply Theorem 3.1, and it remains to show that the resulting sequence of regular conditional probabilities $\mathrm{P}_{X|Y_{n}}$ converges to the regular conditional probability $\mathrm{P}_{X|Y}$. The following definition introduces the type of projections we need for this approach.
@@
The measure-theoretic description of filtering sequences is useful for establishing our general result on conditioning jointly Gaussian random variables, yet it is somewhat cumbersome to verify. The following proposition makes this verification substantially easier.
Proposition 3.3.
Then $(A_{n})$ is a filtering sequence for $F$.
Using Proposition 3.3 it is easy to construct filtering sequences, both in an abstract setting as well as in concrete examples. This will be discussed in detail in Section 4.
@@
Recall that [33] also obtain a nuclear norm convergence for their approximation scheme in the Hilbert space setting. Moreover, note that in the observational model (5), we have ${\cal G}(X,Y)={\cal G}(X)$ and ${\cal G}(Y)\subset{\cal G}(X)$, see Lemma 5.4, and therefore the conditional expectation considered in vi) becomes the orthogonal projection $\Pi_{Y}:{\cal G}(X)\to{\cal G}(X)$ onto the subspace ${\cal G}(Y)$.
Our next goal is to refine our analysis for both $\mathrm{P}_{X|Y}$ and its approximation. To this end, let us assume that we are in the situation of Theorem 3.4. Then each $\mu_{X|Y_{n}}\circ A_{n}:F\to E$ is strongly measurable, where we use the $\sigma$-algebra $\sigma(F^{\prime})=\mathfrak{B}(F)$ on $F$. Let us define $\mu_{X|Y}:F\to E$ by
@@
and since $Z$ is a version of $\mathbb{E}(X|Y)$, we see that $\mu_{X|Y}$ is a version of the “factorized” conditional expectation $\mathbb{E}(X|Y=\,\cdot\,):F\to E$, see also the discussion at the beginning of [27, Chapter 8.3] for the case of $\mathbb{R}$-valued $X$. Consequently, Theorem 3.4 shows that for each version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times\mathbb{R}^{n}\to[0,1]$ we can construct a version $\mu_{X|Y}$ of $\mathbb{E}(X|Y=\,\cdot\,):F\to E$ by setting
@@
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be separable Banach spaces, and $X:\Omega\to E$ and $Y:\Omega\to F$ be jointly Gaussian random variables. Then (17) defines a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$ for all measurable $\mu:F\to E$, for which
@@
is a version of $\mathbb{E}(X|Y)$. In particular, in the case $F=\mathbb{R}^{n}$, these versions satisfy the assertions of Theorem 3.1 for $N=\emptyset$.
@@
Let us assume for a moment, that $X$ and $Y$ are centered. Then $F_{\mathrm{conv},\mathfrak{A}}=F_{X|Y,\mathfrak{A}}$ holds, and therefore $F_{\mathrm{conv},\mathfrak{A}}$ is a measurable subspace of full measure $\mathrm{P}_{Y}$. Moreover, for all $y\in F_{\mathrm{conv},\mathfrak{A}}$ we have
@@
Consequently, the restriction of $\mu_{X|Y,\mathfrak{A}}$ to $F_{\mathrm{conv},\mathfrak{A}}$ is linear and measurable. Note that such linear maps also occurred in [31] when considering conditional Gaussian distributions in the Hilbert space setup.
The theorems above also provide some concentration result. To illustrate this, let us assume that we have an $E$-valued Gaussian random variable $X$ and that we have fixed a filtering sequence $\mathfrak{A}:=(A_{n})$ for $E$. Moreover, assume that we cannot observe $X$ but only $X_{n}:=A_{n}\circ X$. This gives us versions $\mathrm{P}_{X|X_{n}}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probabilities of $X$ given $X_{n}$ with

### v3

@@ 2 Preliminaries
In addition, recall that the $w^{*}$-topology on $E^{\prime}$ is the smallest topology, such that all maps $\langle\,\cdot\,,x\rangle_{E^{\prime},E}:E^{\prime}\to\mathbb{R}$ are continuous. For separable Banach spaces, the $w^{*}$-topology has some highly useful properties, which are summarized in the following result, whose first two assertions can be found in in [34, Theorem 2.6.18 and Theorem 2.6.23] or [13, Theorem V.3.1 and Theorem V.5.1]. The last assertion directly follows from the fact that compact metric spaces are separable.
@@
Note that for the proofs of our main results we need the existence of such sets $\mathfrak{D}$, and another key result, namely Theorem 4.2, also need such sets in its formulation. In addition, the fact that the $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable is often used in the the proofs as it allows to deal with sequences rather than nets and this in turn opens the door for limit theorems of integrals.
In this work, we need two $\sigma$-algebras on $E$: The first one is the usual Borel $\sigma$-algebra $\mathfrak{B}(E)$ that is generated by the norm topology on $E$ and the second one is $\sigma(E^{\prime})$, that is, the smallest $\sigma$-algebra on $E$, for which all bounded, linear $x^{\prime}:E\to\mathbb{R}$ are measurable, where, as usual, $\mathbb{R}$ is equipped with $\mathfrak{B}:=\mathfrak{B}(\mathbb{R})$. We further write $\mathfrak{B}^{d}:=\mathfrak{B}(\mathbb{R}^{d})$ and assume throughout this work that $\mathbb{R}^{d}$ is equipped with $\mathfrak{B}^{d}$.
In general, we have $\sigma(E^{\prime})\subset\mathfrak{B}(E)$ and the following version of Pettis’ measurability theorem, which is a consequence of e.g. [12, Theorem E.9] and which can be also found in [24, Proposition 1.1.1], shows that for separable $E$ both $\sigma$-algebras coincide.
@@
Let $(\Omega,{\cal A})$ be a measurable space, $E$ be a Banach space, and $X:\Omega\to E$. Then $X$ is called weakly measurable if it is $({\cal A},\sigma(E^{\prime}))$-measurable. By a standard measurability argument, see e.g. [29, Corollary 1.82], we see that this is equivalent to the $({\cal A},\mathfrak{B})$-measurability of $\langle x^{\prime},X\rangle_{E^{\prime},E}:\Omega\to\mathbb{R}$ for all $x^{\prime}\in E^{\prime}$. Moreover, $X$ is called strongly measurable if it is $({\cal A},\mathfrak{B}(E))$-measurable and $f(\Omega)$ is separable. If $E$ is separable, then one can quickly check with Theorem 2.2 that weak and strong measurability are equivalent. Finally, if $(\Omega,{\cal A},\mathrm{P})$ is a probability space, then a strongly measurable $X:\Omega\to E$ is called a random variable, and in this case its image measure $\mathrm{P}_{X}$ on $\mathfrak{B}(E)$ is called the distribution of $X$. Of course, if $E$ is separable, every weakly measurable $X:\Omega\to E$ is a random variable, and for this reason we sometimes simply speak of measurable $X$.
@@
It turns out, that in this case the limit $\mathrm{P}$ is unique, see e.g. [50, Chapter I.3.7]. Of course, for $E=\mathbb{R}^{d}$, Lévy’s continuity theorem shows that weak convergence of probability measures can be characterized by the pointwise convergence of the corresponding characteristic functions, see also Theorem C.3. A generalization to the case of separable Banach spaces $E$ is recalled in Theorem C.6. Finally, like in the finite dimensional case, we say that a sequence $(X_{n})$ of $E$-valued random variables converges to an $E$-valued random variable $X$ in distribution, if the distributions of $(X_{n})$ weakly converge to the distribution of $X$.
@@
$\mathbb{E}X:=\int_{\Omega}X\,\mathrm{d}\mathrm{P}\,,$
see also Appendix A, which collects some information on Bochner integrals. Moreover, for $X\in{{\cal L}_{2}(\mathrm{P},E)}$ and $Y\in{{\cal L}_{2}(\mathrm{P},E)}$, where $F$ is another Banach space, the cross covariance of $X$ and $Y$ is the bounded linear operator $\operatorname{cov}(X,Y):F^{\prime}\to E$ defined by the Bochner integrals
@@
Recall that $\operatorname{cov}(X,Y):F^{\prime}\to E$ is a bounded linear operator, which is nuclear and thus compact, see e.g. [50, Theorem 2.5 in Chapter III.2] and Lemma B.7 for an alternative argument. In particular, the covariance of $X$, which is defined by
@@
where $X_{i}$ and $Y_{j}$ denote the $i$th, respectively $j$th, component of $X$ and $Y$. for later use note that in this description, we formally write elements of $(\mathbb{R}^{n})^{\prime}$ as row vectors and identify $(\mathbb{R}^{n})^{\prime}$ by $\mathbb{R}^{n}$ via the canonical isomorphism given by the transposition, that is $\,\cdot\,^{\mathsf{T}}:(\mathbb{R}^{n})^{\prime}\to\mathbb{R}^{n}$. In other words, we have $\operatorname{cov}(X,Y)(y^{\prime})=K_{X,Y}\cdot(y^{\prime})^{\mathsf{T}}$ for all $y^{\prime}\in(\mathbb{R}^{n})^{\prime}$.
@@
Moreover, a probability measure $\mathrm{Q}$ on $\sigma(E^{\prime})=\mathfrak{B}(E)$ is a Gaussian measure, if the image measure $\mathrm{Q}_{x^{\prime}}$ on $(\mathbb{R},\mathfrak{B})$ is a Gaussian measure for all $x^{\prime}\in E^{\prime}$.
It turns out, that it suffices to verify the Gaussianity for a $\mathfrak{D}\subset B_{E^{\prime}}$ that is $w^{*}$-dense in $B_{E^{\prime}}$, see Theorem 6.1. In addition, it directly follows from the definition that $\mathrm{Q}$ is a Gaussian measure on $\sigma(E^{\prime})$, if and only if the identity map $\operatorname{id}_{E}:E\to E$ defined on the probability space $(E,\sigma(E^{\prime}),\mathrm{Q})$ is a Gaussian random variable. Finally, every affine linear, continuous transformation of a Gaussian random variable is a Gaussian random variable. To be more precise, if $X:\Omega\to E$ is a Gaussian random variable, $T:E\to F$ is a bounded linear operator mapping into a separable Banach space $F$, and $y\in F$, then $y+T\circ X$ is Gaussian random variable.
Recall that Fernique’s theorem, see e.g. [32, Corollary 3.2], shows that Gaussian random variables $X:\Omega\to E$ on some probability space $(\Omega,{\cal A},\mathrm{P})$ satisfy
@@
Like in the finite dimensional case, the expectation and covariance uniquely determine a Gaussian measure. Namely, if $\mathrm{P}$ and $\mathrm{Q}$ are Gaussian measures on $\sigma(E^{\prime})$ with $\mu_{\mathrm{P}}=\mu_{\mathrm{Q}}$ and $\operatorname{cov}(\mathrm{P})=\operatorname{cov}(\mathrm{Q})$, then we have $\mathrm{P}=\mathrm{Q}$. This can be quickly verified with the help of characteristic functions, e.g. by combining [50, Theorem 2.2 in Chapter IV.2.1] with [50, Proposition 2.8 in Chapter IV.2.4], where the latter is restated in a simplified form in Proposition C.5. For this reason, we write
@@
for Gaussian measures $\mathrm{Q}$.
Finally, given some Gaussian random variable $X:\Omega\to E$ on a probability space $(\Omega,{\cal A},\mathrm{P})$ we call
@@
the Gaussian Hilbert space associated to $X$, where we identify the functions $\langle x^{\prime},X-\mathbb{E}X\rangle_{E^{\prime},E}:\Omega\to\mathbb{R}$ with their $\mathrm{P}$-equivalence classes. It turns out that ${\cal G}(X)$ is indeed a Gaussian Hilbert space in the sense of [26]. For this and additional properties of ${\cal G}(X)$ we refer to Lemma 6.4.
@@
of $\mathrm{Q}$ given $F$. We write $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):=\mathrm{Q}(\,\cdot\,|\,\cdot\,)$ and speak of the regular conditional probability of $X$ given $Y$. Note that this is justified as $\mathrm{P}(B|y)$ describes the probability of $\{X\in B\}$ under the condition that we have observed $y=Y(\omega)$. Consequently, our goal is to describe $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ modulo $\mathrm{P}_{Y}$-sets. This task will be accomplished by
@@
In the finite dimensional case, the solution of this task is folklore, see e.g. [39, Appendix A.2], [7, Chapter 2.3], [22, Lemma 4.3], [45, Theorem 6.20], and also [38, Section 4]. To formulate the solution in the most general setting, let us recall that an $n\times m$-Matrix $B$ is said to be a Moore-Penrose pseudo-inverse of a given $m\times n$-matrix $A$, if the following four equations hold true:
@@
It can be shown that for every $m\times n$-matrix $A$ there exists exactly one Moore-Penrose pseudo-inverse, which we denote by $A^{\dagger}$, see e.g. [37, Theorem 1]. Similarly, if we have a linear map $A:\mathbb{R}^{n}\to\mathbb{R}^{m}$, then a linear map $B:\mathbb{R}^{m}\to\mathbb{R}^{n}$ is a Moore-Penrose pseudo-inverse of $A$, if the four equations above are satisfied. The uniqueness of such a pseudo-inverse can be shown analogously to the matrix case and the existence follows from the matrix case by using the matrices representing $A$ and $B$. We again write $A^{\dagger}$ for the pseudo-inverse.
@@
Finally, we also need conditional expectations of Banach space valued random variables. To recall the definition, which is a straightforward generalization from the $\mathbb{R}$-valued case, let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, ${\cal C}\subset{\cal A}$ be a sub-$\sigma$-algebra, and $X\in{{\cal L}_{1}(\mathrm{P},E)}$. Then a $\mathrm{P}$-integrable and strongly ${\cal C}$-measurable $Z:\Omega\to E$ is called a conditional expectation of $X$ given ${\cal C}$, if
@@
Like in the $\mathbb{R}$-valued case, the conditional expectation of $X$ given ${\cal C}$ exists and is $\mathrm{P}$-almost surely unique. Hence we write $\mathbb{E}(X|{\cal C}):=Z$ and say that $Z$ is a version of $\mathbb{E}(X|{\cal C})$. Moreover, if ${\cal C}$ is generated by a random variable $Y$, that is ${\cal C}=\sigma(Y)$, we write $\mathbb{E}(X|Y):=\mathbb{E}(X|\sigma(Y))=\mathbb{E}(X|{\cal C})$. Banach space valued conditional expectations enjoy many of the properties of $\mathbb{R}$-valued conditional expectations, see Appendix D. In addition, Appendix D also contains a detailed description of the relationship between regular conditional probabilities and conditional expectations in the Banach space setting.
@@ 3 Main Results
As mentioned in the introduction, [8, Corollary 3.10.3] has already shown that there is a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, such that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian for all $y$. Later, in Theorem 3.4 we will recover such a version, too. Moreover, for more specific cases, formulas for the mean and covariance have also been derived by other authors as laid out in the introduction. However, we could not find such formulas in the full generality presented in Theorem 3.1. Since these formulas play a crucial role for our approximation scheme in the case of infinite dimensional $Y$, we thus decided to present and prove Theorem 3.1 as a result.
Our next and main goal of this section is to generalize Theorem 3.1 to infinite dimensional $Y$. Here we recall that we cannot expect a literal generalization. To illustrate this, let us assume for simplicity that $H:=F$ is an infinite dimensional Hilbert space. Using the canonical Riesz-Fréchet-identification of $H^{\prime}$ with $H$, we can then view $\operatorname{cov}(Y)$ as a compact linear operator $\operatorname{cov}(Y):H\to H$. Now recall that a bounded linear operator $A:H\to H$ has a Moore-Penrose inverse if and only if $\operatorname{ran}A$ is closed, see e.g. [21, Theorem 2.4]. Moreover, it is well-known that compact operators have a closed range, if and only if their range is actually finite dimensional, see e.g. [34, Proposition 3.4.6]. Consequently, the Moore-Penrose inverse $(\operatorname{cov}Y)^{\dagger}$ exists if and only if $\operatorname{ran}\operatorname{cov}(Y)$ is finite dimensional, and the latter is only possible, if the range of $Y$ is almost surely contained in a finite-dimensional subspace of $F$, see e.g. [31, Theorem 2.1] or [46, Theorem 2.43].
In the following, we will resolve this issue by considering a sequence $(Y_{n})$ of suitably chosen finite dimensional “projections” of $Y$. For each $Y_{n}$ we can then apply Theorem 3.1, and it remains to show that the resulting sequence of regular conditional probabilities $\mathrm{P}_{X|Y_{n}}$ converges to the regular conditional probability $\mathrm{P}_{X|Y}$. The following definition introduces the type of projections we need for this approach.
@@
The measure-theoretic description of filtering sequences is useful for establishing our general result on conditioning jointly Gaussian random variables, yet it is somewhat cumbersome to verify. For this reason we present some simple approaches for constructing filtering sequences in Section 4. In addition, we will see there that for separable Banach spaces $F$ there always exists a filtering sequence and that for important classes of spaces filtering sequences can be easily constructed and applied. As an illustrative example, we will work out the details for spaces of continuous function on compact metric spaces in Section 5.
@@
Recall that [36] also obtain a nuclear norm convergence for their approximation scheme in the Hilbert space setting. However, the approximation scheme in [36] is quite distinct from ours. Besides [36] we are not aware of any approximation of $C_{X|Y}$ with respect to the nuclear norm, while known weaker approximations in restricted settings have already been discussed in the introduction.
Moreover, note that in the observational model (5), we have ${\cal G}(X,Y)={\cal G}(X)$ and ${\cal G}(Y)\subset{\cal G}(X)$, as we can quickly derive from Lemma 6.4, and therefore the conditional expectation considered in vi) becomes the orthogonal projection $\Pi_{Y}:{\cal G}(X)\to{\cal G}(X)$ onto the subspace ${\cal G}(Y)$.
Our next goal is to refine our analysis for both $\mathrm{P}_{X|Y}$ and its approximation. To this end, let us assume that we are in the situation of Theorem 3.3. Then each $\mu_{X|Y_{n}}\circ A_{n}:F\to E$ is continuous and thus strongly measurable, where we use the $\sigma$-algebra $\sigma(F^{\prime})=\mathfrak{B}(F)$ on $F$. Let us define $\mu_{X|Y}:F\to E$ by
@@
and since $Z$ is a version of $\mathbb{E}(X|Y)$, we see that $\mu_{X|Y}$ is a version of the “factorized” conditional expectation $\mathbb{E}(X|Y=\,\cdot\,):F\to E$, see also the discussion at the beginning of [29, Chapter 8.3] for the case $E=\mathbb{R}$.
Now assume that we have a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times\mathbb{R}^{n}\to[0,1]$. By Theorem 3.3 we can then alternatively construct a version $\mu_{X|Y}$ of $\mathbb{E}(X|Y=\,\cdot\,):F\to E$ by setting
@@
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be separable Banach spaces, and $X:\Omega\to E$ and $Y:\Omega\to F$ be jointly Gaussian random variables. Moreover, let $\mu:F\to E$ be a measurable such that
@@
is a version of $\mathbb{E}(X|Y)$. Then (15) defines a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$. Moreover, in the case $F=\mathbb{R}^{n}$, these versions satisfy the assertions of Theorem 3.1 for $N=\emptyset$.
@@
Note that if $X$ and $Y$ are centered, then $F_{\mathrm{conv},\mathfrak{A}}=F_{X|Y,\mathfrak{A}}$ holds, and therefore $F_{\mathrm{conv},\mathfrak{A}}$ is a measurable subspace of full measure $\mathrm{P}_{Y}$. Moreover, for all $y\in F_{\mathrm{conv},\mathfrak{A}}$ we have
@@
Consequently, the restriction of $\mu_{X|Y,\mathfrak{A}}$ to $F_{\mathrm{conv},\mathfrak{A}}$ is linear and measurable. Note that such linear maps also occurred in [33], where conditional Gaussian distributions in the Hilbert space case were considered.
Let us finally discuss how the theorems above also provide some frequentist concentration result for the Bayesian method given by the prior distribution $\mathrm{P}_{X}$ and the conditional distributions. To this end, let us assume that we have an $E$-valued Gaussian random variable $X$ and that we have fixed a filtering sequence $\mathfrak{A}:=(A_{n})$ for $E$. Moreover, assume that we cannot observe $X$ but only $X_{n}:=A_{n}\circ X$. This gives us versions $\mathrm{P}_{X|X_{n}}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probabilities of $X$ given $X_{n}$ with